OpenStudy (anonymous):

1)Prove the statement is true 2)Is the converse true? if so, prove it 3)Is the contrapositive true? if so, prove it 4)Is the negation true? if so, prove it

OpenStudy (anonymous):

for all real numbers x and y, if \[x -\lfloor x \rfloor< \frac{ 1 }{ 2} \] and \[y -\lfloor y \rfloor< \frac{ 1 }{ 2} \] then \[\lfloor x + y \rfloor = \lfloor x \rfloor + \lfloor y \rfloor\]

OpenStudy (ikram002p):

[...] define the greatest integer right ?

OpenStudy (anonymous):

umm i dont know

OpenStudy (anonymous):

the statement is above, the question doesnt say find the greatest integer, it asks to prove that the above is true if you assume x and y are real numbers

OpenStudy (usukidoll):

we have an If P then Q statement... and that's all I could give :(

OpenStudy (anonymous):

thanks...all i have is suppose x,y are reals, suppose P is true

OpenStudy (anonymous):

dont know where to start

OpenStudy (usukidoll):

do you know the definitions and axioms?

OpenStudy (ikram002p):

ok its like this if u have y=3.18.. then [y]=4 ok ?

OpenStudy (usukidoll):

these proof crud things are like well you need to apply the right definition and then think outside the dam n box

OpenStudy (usukidoll):

I'm studying it now, but my class is doing induction

OpenStudy (usukidoll):

just finished the Tower of Hanoi

OpenStudy (anonymous):

yes i have definitions for floor and ceilings, but dunno how i would use them, it would introduce more variables and etc

OpenStudy (ikram002p):

x-[x]<0.5 is true y-[y]<0.5 is true proof [x+y]=[x]+[y]

OpenStudy (ikram002p):

brb ok ?

OpenStudy (anonymous):

k

OpenStudy (usukidoll):

the negation would be everything negative on this statement

OpenStudy (anonymous):

i just know my instructor said something about adding the left side of the inequalities together

OpenStudy (usukidoll):

the negation is those two equations you got are greater than 1/2

OpenStudy (usukidoll):

and the end result is that they shouldn't equal each other

OpenStudy (anonymous):

x−⌊x⌋ + y−⌊y⌋ < 1...i think

OpenStudy (usukidoll):

O_O

OpenStudy (anonymous):

x + y < 1 + ⌊x⌋ + ⌊y⌋

OpenStudy (anonymous):

i dunno, lol

OpenStudy (usukidoll):

:O no :_:

ganeshie8 (ganeshie8):

add first two given conditions

OpenStudy (anonymous):

1 + ⌊x⌋ + ⌊y⌋ is the ceiling of x+y?

ganeshie8 (ganeshie8):

\(x -\lfloor x \rfloor< \frac{ 1 }{ 2} \) \(y -\lfloor y \rfloor< \frac{ 1 }{ 2} \)

OpenStudy (anonymous):

and thus it's floor is ⌊x⌋ + ⌊y⌋ ..

ganeshie8 (ganeshie8):

adding, \(\large x + y - \lfloor x \rfloor - \lfloor y \rfloor < 1 \)

ganeshie8 (ganeshie8):

\(\large x + y < 1 + \lfloor x \rfloor + \lfloor y \rfloor \)

ganeshie8 (ganeshie8):

now take the floor both sides

ganeshie8 (ganeshie8):

\(\large \lfloor x + y \rfloor < \lfloor 1 + \lfloor x \rfloor + \lfloor y \rfloor \rfloor \)

OpenStudy (anonymous):

\[\lfloor x \rfloor + \lfloor y \rfloor \le x + y < \lfloor x \rfloor + \lfloor y \rfloor +1\]

ganeshie8 (ganeshie8):

floor of x and floor of y gives u the sum of integer values, and floor of a number less than 1 is simply 0

ganeshie8 (ganeshie8):

so, \(\large \lfloor x + y \rfloor = \lfloor x \rfloor + \lfloor y \rfloor \)

ganeshie8 (ganeshie8):

so the given statement is true

OpenStudy (anonymous):

can you explain the right side again please?

ganeshie8 (ganeshie8):

\(\large \lfloor x + y \rfloor < \lfloor 1 + \lfloor x \rfloor + \lfloor y \rfloor \rfloor \) take for example : x = 3.4, y = 7.2 [x] = 3 [y] = 7 [<1] = 0

ganeshie8 (ganeshie8):

right hand evaluates to 0+3+7 = 10, which is same as left hand side : [3.4 + 7.2] = [10.6] = 10

ganeshie8 (ganeshie8):

that example is oly for you to *see* thats not required for the main proof..

OpenStudy (anonymous):

yeah, im trying to show this without subbing in numbers

ganeshie8 (ganeshie8):

'was already shown before the example.. go thru again :P

OpenStudy (anonymous):

⌊1+⌊x⌋+⌊y⌋⌋ to ⌊x⌋+⌊y⌋ im still confused about this

ganeshie8 (ganeshie8):

if u digest below, it wud make sense :- if 0 < a < 1, then [a] = 0

ganeshie8 (ganeshie8):

[0.99] = [0.1] = [0.4] = 0

ganeshie8 (ganeshie8):

see if below argument makes sense :- say x-[x] = a and y-[y] = b then from the hypothesis, a+b < 1 adding the given statements : x+y = a+b + [x] + [y] taking floor both sides : [x+y] = [ a+b + [x] + [y]] = [(a+b) + integer] = [a+b] + integer = 0 + integer = [x] + [y]